# Coupled Dipole Oscillation Enabled Scintillation Compensation in Turbulence-Impacted Laser Beam

### v1

@@ 1 Introduction
Optical communication systems operating in real-world environments often face severe performance degradation due to atmospheric turbulence, which causes random fluctuations in the intensity, phase, and polarization of propagating beams. This results in beam wander, scintillation, and loss of coherence, ultimately reducing the reliability of free-space optical links. The work presented in this study introduces a rigorous theoretical and experimental framework for mitigating such turbulence-induced distortions by leveraging collective dipole oscillations in dense dielectric media.
@@
This work presents a comprehensive range of studies that collectively advance the fundamental understanding of optical forces, electromagnetic dipoles, and wave-matter interactions. Beginning with mathematical and statistical frameworks, Vishwakarma and Moosath propose innovative methods for quantifying distances in the space of Gaussian mixtures, pivotal for signal analysis and probability theory [1]. Early studies by Korotkova and collaborators systematically examine the evolution of intensity fluctuations and Stokes parameter statistics of random and quasi-monochromatic electromagnetic beams in turbulent or complex media, providing extensive insights on coherence, polarization, and the role of environmental randomness [2, 3, 4, 5, 6, 7, 8, 9, 10]. The efforts to forge new calculi for optical angular momentum transformations via Jones and Stokes-Mueller matrices have also expanded the mathematical toolkit for polarized beam diagnostics [11, 12, 13, 14, 15]. Abbasirad et al.’s investigation into dipole emission near dielectric metasurfaces with scanning near-field optical microscopy underscores advanced experimental techniques for probing electromagnetic responses at the nanoscale [16], while Scheel and Buhmann’s treatise on macroscopic quantum electrodynamics broadens the conceptual base for light-matter coupling in engineered and natural settings [17]. Classic works such as Levine and Schwinger’s aperture diffraction theory and Sipe’s macroscopic electromagnetic description of resonant dielectrics have shaped the theoretical foundations that underpin contemporary treatments of optical scattering and absorption phenomena [18, 19].
@@
Multiple contributions address the dynamics, enhancement, and nonlinearities of dipole-dipole interactions as well as the ramifications for metamaterials and nanocomposites. Poddubny et al. propose microscopic models for Purcell enhancement within hyperbolic metamaterials, crucial for light emission control [20]. Martin and Piller, as well as Paulus et al., deliver accurate computational techniques for electromagnetic scattering and Green’s tensor evaluation in complex stratified environments [21, 22]. The evolving understanding of the Lorentz oscillator, including nonlinear effects and the ramifications of radiative damping, is enriched by Smith and Schelew et al. [23, 29]. Studies on optical forces—including fundamental reviews, force and torque predictions by spinning fields, and experimental observations of light-induced dipole forces in ultracold gases—are presented by Tamura [33], Lembessis [34], Saif [35], Matsumori [37], Canaguier-Durand [38], Kumar [39], and Maiwoger [40], demonstrating the breadth and technological relevance of optomechanical manipulation. Further, several references deepen the discourse on discrete dipole approximations, time-varying polarizabilities, quadrupolar responses, and nonlinearity in nanoantennas and metasurfaces, such as Bowen [25], Mirmoosa [26], Achouri [30], and Sain [45]. Collectively, these papers establish a rigorous, multidisciplinary framework that informs the present investigation and inspires future advances in optical physics, material science, and wave-based technologies.
@@
The theoretical foundation begins with the Lorentz oscillator model, which describes the oscillation of bound electrons under an external electromagnetic field. While the classical model accounts for linear restoring forces, real molecular systems—such as those in polymer media—also exhibit nonlinear restoring contributions. To address this, the model is extended to include both quadratic (second-order) and cubic (third-order) nonlinearities. These anharmonic corrections capture more realistic electron cloud dynamics and their nonlinear interaction with incident fields. A central theme of the work is the role of dipole–dipole interactions, which are introduced through the dyadic Green’s function formalism. Unlike the scalar Green’s function, the dyadic representation preserves the vector nature of electromagnetic fields and accurately accounts for near-field, induction-zone, and far-field coupling between oscillating dipoles. This nonlocal description of the medium enables the derivation of collective oscillation modes, obtained by diagonalizing the effective stiffness matrix of the system. The resulting eigenmodes form a natural basis for perturbative analysis: zeroth-order solutions describe the linear response, while higher orders incorporate nonlinear corrections that manifest in effective nonlocal susceptibility kernels. The output field is reconstructed using the scalar Green’s function for free-space propagation, linking the microscopic dipole oscillations to macroscopic beam statistics. This theoretical framework predicts that when light propagates through a dielectric medium—such as PMMA (poly(methyl methacrylate)) rods—the collective synchronization of dipole oscillations can significantly reduce scintillation caused by turbulence. Physically, synchronized dipoles impose inertia on fluctuating external fields, slowing down centroid shifts and suppressing intensity variance.
@@
When an electromagnetic field propagates through a transparent or semi-transparent medium, its electric field interacts with the molecular structures of the material. This interaction perturbs the electron clouds, breaking the stationary symmetries of the molecules and giving rise to induced dipoles. As the external field oscillates, it drives corresponding dipole oscillations within the medium. Since the molecules are not isolated but interconnected, they are able to exchange energy, leading to the emergence of a coupled dipole oscillation model. In such a system, the coupling imposes constraints on the motion of individual dipoles, effectively altering the system’s degrees of freedom. The coupling coefficients introduce non-diagonal terms into the governing equations, meaning the standard Cartesian coordinates can no longer serve as an orthonormal basis for describing the dynamics. Consequently, the system cannot be represented by independent dynamical equations without further transformation. To resolve this, a process of diagonalization is applied to the coupled equations, redefining the degrees of freedom and establishing a new orthonormal reference frame in which the system can be analyzed more clearly. The diagonalized representation reveals a set of collective oscillatory modes that describe the synchronized dipole behavior. Unlike the original random dipole oscillations, which are chaotic and uncorrelated, the coupled mode dynamics reflect coherent, collective oscillations stabilized by the mutual interactions among dipoles. This transformation from random to coupled modes highlights the importance of diagonalization in simplifying the description of complex dipole systems, allowing the emergent ordered dynamics to be distinguished from the initial disorder induced by the external field. Ultimately, this framework provides a powerful way to understand how microscopic molecular interactions shape the macroscopic electromagnetic response of materials.

### v2

@@ 1 Introduction
Optical communication systems operating in real-world environments often face severe performance degradation due to atmospheric turbulence, which causes random fluctuations in the intensity, phase, and polarization of propagating beams. This results in beam wander, scintillation, and loss of coherence, ultimately reducing the reliability of free-space optical links. The work presented in this study introduces a rigorous theoretical and experimental framework for mitigating such turbulence-induced distortions by leveraging collective dipole oscillations in dense dielectric media.[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]
@@
This work presents a comprehensive range of studies that collectively advance the fundamental understanding of optical forces, electromagnetic dipoles, and wave-matter interactions. Beginning with mathematical and statistical frameworks, Vishwakarma and Moosath propose innovative methods for quantifying distances in the space of Gaussian mixtures, pivotal for signal analysis and probability theory [1]. Early studies by Korotkova and collaborators systematically examine the evolution of intensity fluctuations and Stokes parameter statistics of random and quasi-monochromatic electromagnetic beams in turbulent or complex media, providing extensive insights on coherence, polarization, and the role of environmental randomness [2, 3, 4, 5, 6, 7, 8, 9, 10]. The efforts to forge new calculi for optical angular momentum transformations via Jones and Stokes-Mueller matrices have also expanded the mathematical toolkit for polarized beam diagnostics [11, 12, 13, 14, 15]. Abbasirad et al.’s investigation into dipole emission near dielectric metasurfaces with scanning near-field optical microscopy underscores advanced experimental techniques for probing electromagnetic responses at the nanoscale [16], while Scheel and Buhmann’s treatise on macroscopic quantum electrodynamics broadens the conceptual base for light-matter coupling in engineered and natural settings [17]. Classic works such as Levine and Schwinger’s aperture diffraction theory and Sipe’s macroscopic electromagnetic description of resonant dielectrics have shaped the theoretical foundations that underpin contemporary treatments of optical scattering and absorption phenomena [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@
Multiple contributions address the dynamics, enhancement, and nonlinearities of dipole-dipole interactions as well as the ramifications for metamaterials and nanocomposites. Poddubny et al. propose microscopic models for Purcell enhancement within hyperbolic metamaterials, crucial for light emission control [20]. Martin and Piller, as well as Paulus et al., deliver accurate computational techniques for electromagnetic scattering and Green’s tensor evaluation in complex stratified environments [21, 22]. The evolving understanding of the Lorentz oscillator, including nonlinear effects and the ramifications of radiative damping, is enriched by Smith and Schelew et al. [23, 29]. Studies on optical forces—including fundamental reviews, force and torque predictions by spinning fields, and experimental observations of light-induced dipole forces in ultracold gases—are presented by Tamura [33], Lembessis [34], Saif [35], Matsumori [37], Canaguier-Durand [38], Kumar [39], and Maiwoger [40], demonstrating the breadth and technological relevance of optomechanical manipulation. Further, several references deepen the discourse on discrete dipole approximations, time-varying polarizabilities, quadrupolar responses, and nonlinearity in nanoantennas and metasurfaces, such as Bowen [25], Mirmoosa [26], Achouri [30], and Sain [45]. Collectively, these papers establish a rigorous, multidisciplinary framework that informs the present investigation and inspires future advances in optical physics, material science, and wave-based technologies.[31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]
@@
The theoretical foundation begins with the Lorentz oscillator model, which describes the oscillation of bound electrons under an external electromagnetic field. While the classical model accounts for linear restoring forces, real molecular systems—such as those in polymer media—also exhibit nonlinear restoring contributions. To address this, the model is extended to include both quadratic (second-order) and cubic (third-order) nonlinearities. These anharmonic corrections capture more realistic electron cloud dynamics and their nonlinear interaction with incident fields. A central theme of the work is the role of dipole–dipole interactions, which are introduced through the dyadic Green’s function formalism. Unlike the scalar Green’s function, the dyadic representation preserves the vector nature of electromagnetic fields and accurately accounts for near-field, induction-zone, and far-field coupling between oscillating dipoles. This nonlocal description of the medium enables the derivation of collective oscillation modes, obtained by diagonalizing the effective stiffness matrix of the system. The resulting eigenmodes form a natural basis for perturbative analysis: zeroth-order solutions describe the linear response, while higher orders incorporate nonlinear corrections that manifest in effective nonlocal susceptibility kernels. The output field is reconstructed using the scalar Green’s function for free-space propagation, linking the microscopic dipole oscillations to macroscopic beam statistics. This theoretical framework predicts that when light propagates through a dielectric medium—such as PMMA (poly(methyl methacrylate)) rods—the collective synchronization of dipole oscillations can significantly reduce scintillation caused by turbulence. Physically, synchronized dipoles impose inertia on fluctuating external fields, slowing down centroid shifts and suppressing intensity variance.[46, 47, 48, 49, 50, 51, 52, 53, 54, 55]
@@
When an electromagnetic field propagates through a transparent or semi-transparent medium, its electric field interacts with the molecular structures of the material. This interaction perturbs the electron clouds, breaking the stationary symmetries of the molecules and giving rise to induced dipoles. As the external field oscillates, it drives corresponding dipole oscillations within the medium. Since the molecules are not isolated but interconnected, they are able to exchange energy, leading to the emergence of a coupled dipole oscillation model. In such a system, the coupling imposes constraints on the motion of individual dipoles, effectively altering the system’s degrees of freedom. The coupling coefficients introduce non-diagonal terms into the governing equations, meaning the standard Cartesian coordinates can no longer serve as an orthonormal basis for describing the dynamics. Consequently, the system cannot be represented by independent dynamical equations without further transformation. To resolve this, a process of diagonalization is applied to the coupled equations, redefining the degrees of freedom and establishing a new orthonormal reference frame in which the system can be analyzed more clearly. The diagonalized representation reveals a set of collective oscillatory modes that describe the synchronized dipole behavior. Unlike the original random dipole oscillations, which are chaotic and uncorrelated, the coupled mode dynamics reflect coherent, collective oscillations stabilized by the mutual interactions among dipoles. This transformation from random to coupled modes highlights the importance of diagonalization in simplifying the description of complex dipole systems, allowing the emergent ordered dynamics to be distinguished from the initial disorder induced by the external field. Ultimately, this framework provides a powerful way to understand how microscopic molecular interactions shape the macroscopic electromagnetic response of materials.[56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]
@@ 4 Experimental Varifications
Classical turbulence arises from irregular velocity fluctuations in viscous fluids such as the atmosphere, where flow can exist in either laminar or turbulent states. Laminar flow is smooth and orderly, while turbulence is dominated by random subflows, or eddies, that enhance mixing. The transition between these regimes is governed by the Reynolds number, $Re=Vl/\nu$, where $V$ is velocity, $l$ the flow scale, and $\nu$ the kinematic viscosity. When $Re$ exceeds a critical value (typically $\sim 10^{5}$ near the ground), turbulence develops. Kolmogorov’s theory describes turbulence as statistically homogeneous and isotropic at small scales, with large-scale energy generated by shear or convection cascading down to smaller eddies. Energy transfer occurs across an inertial range bounded by an outer scale $L_{0}$ and an inner scale $l_{0}$, until dissipation converts residual energy into heat. The corresponding turbulence power spectrum is expressed as
$\Phi(k)=0.023r_{0}^{-5/3}k^{-11/3}.$ (90)
The Pseudo Random Phase Plate (PRPP) used in our experiment is a five-layered optical device designed to replicate atmospheric turbulence. It consists of two BK7 glass windows enclosing a central acrylic layer imprinted with a Kolmogorov-type turbulence profile. Near-index-matching polymer layers on each side of the acrylic provide mechanical stability, while the glass sealing enhances durability and minimizes environmental effects. The plate, about 10 mm thick, is robust and easily mounted on a rotary stage. It generates aberrated wavefronts with adjustable Fried coherence lengths ($r_{0}$ = 16–32 samples) across 4096 phase points, enabling controlled turbulence simulations.